# Hypergraph Representations of scRNA-seq Data for Improved Clustering with Random Walks

### v3

@@ III Experiments and Results
III.1 Baseline Methods
To evaluate the performance of DIPHW and CoMem-DIPHW, we compared them to several widely used and established cell clustering methods, including five embedding-based methods (PCA [43], Node2Vec [44], EDVW [41], UMAP, and t-SNE) and four unipartite-graph-based community detection methods (Louvain, Greedy Modularity, Infomap, and Leiden). These methods are adopted by popular scRNA-seq analysis packages such as Seurat [15], WGCNA [14], and Scanpy [17]. We provide a summary of these methods in Table 2. For a fair comparison under consistent conditions, we focus on the core clustering algorithms rather than specific implementations in packages such as Seurat and Scanpy. This allows us to evaluate the clustering capabilities of each method independently of any additional preprocessing or optimization steps. The hyperparameters used for each experiment are detailed in Table S1 in the supplementary materials Supplementary Materials.
Method Type Data Input Packages
Greedy
Modularity [45]
Community Detection Unipartite network
(Cell co-expression network)
igraph [46]
Louvain [47] Community Detection Unipartite network
(Cell co-expression network)
Seurat [15], Scanpy [17], WGCNA [14]
Infomap [48] Community Detection Unipartite network
(Cell co-expression network)
igraph [46]
Leiden [49] Community Detection Unipartite network
(Cell co-expression network)
igraph [46]
PCA [43] Embedding Matrix (Genes by cells) Seurat [15], Scanpy [17], SC3 [50], Cell Ranger [26]
t-SNE [51] Embedding Matrix (Genes by cells) Seurat [15], Scanpy [17]
Node2Vec [44] Embedding Unipartite network
(Cell co-expression network)
Multiple implementations available [52, 53] .
UMAP [54] Embedding Matrix (Genes by cells) Seurat [15], Scanpy [17]
EDVW [41] +
word2vec [55]
Embedding Hypergraph incidence matrix
(Genes by cells)
Implemented based on Chitra and Raphael [41]; available in our GitHub repository [56]. Not in standard scRNA-seq packages.
DIPHW +
word2vec [55]
Embedding Hypergraph incidence matrix
(Genes by cells)
Our proposed method and implementation is available on GitHub [56]
CoMem +
word2vec [55]
Embedding Hypergraph incidence matrix
(Genes by cells);
Gene co-expression network;
Cell co-expression network
Our proposed method and implementation is available on GitHub [56]
Table 2: Comparison of Clustering Methods Evaluated in Experiments. For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
@@ III.2 Data and Preprocessing
The data used in this study consists of both simulated and publicly available scRNA-seq datasets. The simulated data were generated using our proposed simulation algorithm described in detail in the Methods section, with code available on GitHub [56]. For preprocessing, we first normalized the raw single-cell RNA-seq data to account for differences in sequencing depth. Genes and cells with zero expression were then removed, and only the top n most variable genes were retained for analysis. For empirical analysis, we used four publicly available scRNA-seq datasets: human brain [57], human pancreas [58], mouse brain [59], and mouse pancreas [60], and a curated benchmarking dataset with ground-truth cell line labels [61].
III.3 Sparsity-Induced Correlation Inflation in scRNA-seq Co-expression Networks
@@
The zero inflation problem in single-cell RNA sequencing (scRNA-seq) data analysis refers to the excessive number of zero counts (or "dropouts") observed in the data. These zeros could either be a result of technical limitation in detecting low abundance transcripts (false negatives), or true absence of biological zeros (true negatives) [37, 39, 40]. The process of integrating individual cell expression profiles into a unified scRNA-seq dataset requires constructing a matrix that accounts for every gene detected in any cell, which artificially introduces zero entries wherever genes are not expressed in those cells. Since cells typically express different sets of genes, this inclusion introduces significant sparsity into the merged dataset. In co-expression network construction, edge weights are inflated due to these artificially introduced common zeros since the common zeros will be considered as evidence of their expression homogeneity. This issue is particularly pronounced at the single-cell level compared to bulk tissue RNAseq data, as the proportion of expressed genes at the single-cell level is much smaller, resulting in higher induced data sparsity.
@@
In the experiment shown in Fig. 3, we investigate how induced sparsity affects correlation coefficients and the interpretation of gene and cell connections in co-expression networks. To simulate this effect, we generated gene expression data with a fixed base size and introduced varying proportions of zeros to mimic different sparsity levels. Our results show that as the level of induced sparsity increases, correlation coefficients between expression profiles become artificially inflated, leading to an overestimation of similarity between cell expression profiles. These findings highlight the importance of addressing zero inflation in scRNA-seq data analysis and the need for accurate representation of the underlying biological relationships. While induced sparsity inflates correlations, our supplementary experiments (.3) show that ignoring shared zeros using cosine similarity leads to worse clustering performance, as these inactive genes still hold biological significance.
[…]
III.4 Simulation Analysis Results
In our study, we developed an algorithm to simulate data that mimics the characteristics of scRNA-seq expression data. In addition to the underlying assumption of within-cell-type homogeneous expression that governs most, if not all, cell clustering methods, we also incorporated between-cell-type crosstalk [62, 63, 64, 65, 66] by integrating configurable intermodular covariance, density, and signal strength into the simulation model. The algorithm generates a sparse matrix representing the simulated scRNA-seq expression data, with numerous user-configurable parameters, including number of modules, module density, shape, background signal strength, modular signal strength, intermodular signal strength, intermodular covariance, and noise level. This method is implemented as a function in our codebase and is available for use by other researchers.
Our simulation model is designed with high flexibility to accommodate different types of scRNA-seq data. For instance, the model can simulate both deeply sequenced datasets with high density (up to 50%) and shallowly sequenced datasets with low density (as low as 1%) [67]. The proportion of non-housekeeping genes can be adjusted by changing the number of differentially expressed genes in each module. Moreover, the variance in module sizes and the number of co-expressed genes in each cell type can be configured by setting different variances for the Poisson distribution used in the embedded module simulation. The modules embedded in the simulated data can be configured with higher density and average signal strength to model the expression profiles of biomarker genes that are highly expressed in certain cell types. This allows us to mimic the sparsity and modularity of real scRNA-seq data, where a small number of genes are highly expressed in specific cell types. Alternatively, the embedding modules can also be user-specified without using the embedded module simulation.
III.4.1 Evaluation Metrics
@@
To evaluate the performance of the clustering methods on datasets with known ground-truth, we employed two widely used metrics for measuring coherence between the identified clusters and the ground-truth clusters: Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI). Both metrics are invariant to permutations, with a score of 0 representing random labeling and a score of 1 indicating identical clusters. For embedding-based methods that do not directly assign cluster membership, we used K-means to partition the embedded data into clusters.
[…]
III.4.2 Clustering Performance vs. Bipartite Modularity and Method Comparison
@@
We evaluated the performance of these clustering methods under varying data modularity conditions. The modularity variation was controlled by two key parameters: the average number of co-expressed genes per cell type and the number of embedded cell modules. The number of modules here represents the number of cell types in the simulated scRNA-seq data. Weak modularity corresponds to a small number of co-expressed genes per module or a large number of embedded modules, making data structures more challenging to detect. The modularity bar plots in Fig. 4 and heatmap visualization in Fig. S6 demonstrate that the simulated scRNA-seq data displays stronger modularity when the average number of co-expressed genes per cell type is greater or when the number of modules is smaller. We implemented Barber’s bipartite modularity [42], as shown by the barplots in Fig. 4, to help quantify modularity of the underlying data for clustering and analyze the association between modularity and clustering performance. Clustering performance generally improves with increased data modularity. Our proposed methods—DIPHW, CoMem, and CoMem-DIPHW—consistently outperformed all other clustering methods that use unipartite network projections, including Node2Vec, and community detection-based methods such as Louvain and Greedy Modularity. Notably, CoMem-DIPHW consistently demonstrated superior performance across almost all parameter settings, with its advantage being particularly pronounced under weak data modularity conditions. For clarity of presentation, we present the most contrasting cases from our analysis: (i) an average of 10 and 100 co-expressed genes per module in Fig. 5 and (ii) 10 and 100 embedded modules in Fig. 6.
[…]
@@
In Fig. 5 and Fig. 6, we compare the performance of our proposed hypergraph-based clustering methods (highlighted in red and yellow) against five embedding-based methods and four community detection methods under both strong and weak modularity conditions. All embedding-based methods are coupled with K-means to produce clustering results. In weak modularity regimes (where the average number of co-expressed genes = 10 or when number of modules = 100), CoMem-DIPHW and DIPHW exhibited a significant performance advantage—for example, when the average number of co-expressed genes = 10, CoMem-DIPHW had an ARI of 0.29 compared to 0.05 for the best-performing baseline method (PCA), representing a performance improvement of approximately 450%. Under strong modularity conditions (when the average number of genes per module = 100 or when number of modules = 10), several methods—including PCA, CoMem-DIPHW, CoMem, Node2Vec, and DIPHW—performed well, achieving average ARI scores above 0.9.
@@
Additional results from intermediate parameter settings (average number of co-expressed genes per module: 20, 30, 40, 70, 80, 90; number of embedded modules: 20, 30, 40, 70, 80, 90) and by NMI support the same observation. Out of the 16 tested scenarios, CoMem-DIPHW or DIPHW consistently ranked first by ARI and in 10 scenarios by NMI. Our hypergraph-based methods showed significant advantages in weak modularity regimes, where data structures are harder to detect. Under comparatively strong modularity conditions, several methods (PCA, CoMem-DIPHW, CoMem, Node2Vec, DIPHW, and Louvain) performed well, with average NMI scores above 0.9. Notably, PCA slightly outperformed CoMem-DIPHW by NMI under strong modularity, with differences in NMI scores ranging from 0.01 to 0.02. These additional results are presented in Supplementary Figs. S2, Fig. S3, Fig. S4, Fig. S5.
III.5 scMixology Benchmark Results
@@
A major challenge in scRNA-seq analysis is the lack of reliable ground-truth for cell type annotation, which complicates the evaluation of clustering algorithms. In most scRNA-seq datasets, cell type labels are typically assigned using a combination of clustering results, differential expression (DE) analysis, and canonical marker genes. First, cells are grouped into clusters based on transcriptomic similarity using clustering algorithms (e.g., Leiden, Louvain, or k-means). Then, DE analysis is performed to identify genes that are differentially expressed across clusters. Finally, these clusters are annotated based on known canonical marker genes associated with specific cell types. Since this annotation process relies on clustering results, using these labels to evaluate clustering performance introduces circular logic and bias toward the method originally used for annotation. To address this issue, we utilized the mixture control benchmarking dataset [61], where ground truth is established based on cell line identity.
@@
We evaluated the performance of our proposed method, CoMem-DIPHW, using two benchmarking datasets from scMixology [61]. These datasets provide ground-truth labels for cell type annotation based on human lung adenocarcinoma cell lines. As shown in Table 3, CoMem-DIPHW outperformed other methods, particularly on the 5-class dataset, achieving the highest NMI and ARI scores. In comparison, Leiden showed strong performance on the 3-class dataset but underperformed significantly on the 5-class dataset. Among the methods evaluated, PCA, CoMem, Node2Vec and CoMem-DIPHW consistently demonstrated robust performance across both datasets.

### v4

@@ III Experiments and Results
III.1 Baseline and Competing Methods
To evaluate the performance of DIPHW and CoMem-DIPHW, we compared them with 13 cell-clustering methods from various categories, including community detection, embedding-based methods, and recent deep learning methods that cluster scRNA-seq data. Table 2 summarizes these methods. For a fair comparison under consistent conditions, we focus on core clustering algorithms rather than specific implementations in packages such as Seurat [satija2015spatial], Scanpy [wolf2018scanpy], and WGCNA [langfelder2008wgcna]. This approach allows us to evaluate the clustering capabilities of each method independently of any additional preprocessing or optimization steps. The hyperparameters used for each experiment are detailed in Table S4 in the supplementary materials.
Method Type Input Packages
Greedy Modularity [newman2004fast] Community detection Cell co-expression network igraph [csardi2006igraph]
Louvain [blondel2008fast] Community detection Cell co-expression network Seurat [satija2015spatial], Scanpy [wolf2018scanpy], WGCNA [langfelder2008wgcna]
Infomap [rosvall2008maps] Community detection Cell co-expression network igraph [csardi2006igraph]
Leiden [traag2019louvain] Community detection Cell co-expression network igraph [csardi2006igraph]
PCA [pearson1901liii] Linear embedding Gene $\times$ cell matrix Seurat [satija2015spatial], Scanpy [wolf2018scanpy], SC3 [kiselev2017sc3], Cell Ranger [zheng2017massively]
t-SNE [van2008visualizing] Manifold embedding Gene $\times$ cell matrix Seurat [satija2015spatial], Scanpy [wolf2018scanpy]
UMAP [mcinnes2018umap] Manifold embedding Gene $\times$ cell matrix Seurat [satija2015spatial], Scanpy [wolf2018scanpy]
Node2Vec [grover2016Node2Vec] Graph embedding Cell co-expression network Multiple implementations available [grover_Node2Vec_2016, snap_Node2Vec_nodate]
graph-sc [ciortan2022gnn] Deep graph embedding Cell co-expression network GitHub [graphsc_github]
tsImpute [zheng2023tsimpute] Imputation Gene $\times$ cell matrix GitHub [tsimpute_github]
CAKE [liu2024cake] Deep contrastive clustering Gene $\times$ cell matrix GitHub [cake_github]
scASDC [min2024scasdc] Deep clustering Gene $\times$ cell matrix GitHub [scasdc_github]
EDVW [Chitra2019] + word2vec [mikolov2013efficient] Hypergraph embedding Gene $\times$ cell matrix Implemented based on [Chitra2019]; GitHub [wan2025comem]
DIPHW + word2vec [mikolov2013efficient] Hypergraph embedding Gene $\times$ cell matrix Our proposed method; GitHub [wan2025comem]
CoMem + word2vec [mikolov2013efficient] Hypergraph embedding Gene $\times$ cell matrix, plus gene and cell co-expression networks Our proposed method; GitHub [wan2025comem]
Table 2: List of Clustering Methods Evaluated in Our Study. K-means was used to cluster the output of all embedding-based methods that do not include a clustering algorithm. Since tsImpute is an imputation method, we evaluated its impact on clustering performance by applying PCA followed by K-means to its imputed data.
@@ III.2 Data and Preprocessing
We use simulated and publicly available scRNA-seq datasets for our study. The simulated data is generated using our simulation algorithm described in Section II.5, with pseudocode in Algorithm 1 and code at [wan2025comem].
For both simulated and real scRNA-seq data, preprocessing includes Counts Per Million normalization [vallejos2017normalizing] to account for differences in sequencing depth, log transformation to reduce skewness, and removal of genes and cells with zero total expression. For real scRNA-seq datasets, we additionally retain the top $n$ highly variable genes using Scanpy’s implementation of the Seurat [satija2015spatial] method. The values of $n$ used for each dataset are provided in Table S4.
We use six publicly available scRNA-seq datasets: (1) human brain [camp2015human], (2) human pancreas [muraro2016single], (3) mouse brain [zeisel2015cell], (4) mouse pancreas [baron2016single], and (5) scMixology benchmark with 3 classes and (6) scMixology benchmark with 5 classes. scMixology benchmark which is a curated benchmark dataset with ground-truth cell line labels [tian2019benchmarking]. Table 2 provides a summary of these datasets.
Datasets  # of Cells  # of Genes  Sparsity Results in
scMixology Benchmark (3-class) [tian2019benchmarking] 902 16,468 45.02% Table 4
scMixology Benchmark (5-class) [tian2019benchmarking] 3,918 11,786 63.01% Table 4
Human Pancreas  [muraro2016single] 3,072 18,348 78.74% Section III.6
Human Brain  [camp2015human] 735 18,929 80.11% Section S7
Mouse Brain  [zeisel2015cell] 3,006 19,973 81.21% Section S8
Mouse Pancreas  [baron2016single] 1,065 14,881 87.79% Section S8
Table 3: List of Publicly Available Datasets Used in Our Study. Sparsity measures the percentage of zero entries in a cell $\times$ gene matrix.
III.3 Sparsity-Induced Inflated Correlation in scRNA-seq Co-expression Networks
@@
The zero inflation problem in scRNA-seq data analysis refers to the excessive number of zero counts (or "dropouts") observed in the data. These zeros may arise from technical limitations in detecting low-abundance transcripts (false negatives) or from the true absence of transcripts (true negatives) [jiang2022statistics, qiu2020embracing, kim2020demystifying]. Integrating individual cell expression profiles into a unified scRNA-seq dataset requires constructing a matrix that includes every gene detected in any cell, which introduces zero entries wherever genes are not expressed in those cells. Because cells typically express different sets of genes, this process creates significant sparsity in the merged dataset. In co-expression network construction, edge weights are inflated due to these artificially introduced common zeros, as the common zeros are interpreted as evidence of expression homogeneity. This issue is especially pronounced at the single-cell level compared to bulk tissue RNA-seq data, since the proportion of expressed genes per cell is much smaller, resulting in greater data sparsity.
@@
In Fig. 3, we examine how induced sparsity affects correlation coefficients and the interpretation of gene and cell connections in co-expression networks. To simulate this effect, we generate gene expression data with a fixed base size and introduce varying proportions of zeros to represent different sparsity levels. Our results indicate that as induced sparsity increases, correlation coefficients between expression profiles become artificially inflated, leading to an overestimation of similarity between cell expression profiles. These findings highlight the importance of addressing zero inflation in scRNA-seq data analysis and the need for accurate representation of underlying biological relationships.
Table S1 (in the Supplementary Section S2) shows that ignoring shared zeros using cosine similarity results in poorer clustering performance, as these inactive genes still have biological significance.
Figure 3: Impact of Increased Sparsity on Correlation Coefficients. This figure shows the relationship between the proportion of zeros added to the expression profile (representing increased sparsity) and the resulting correlation coefficients. This process was repeated 100 times. Shaded areas indicate the 95% confidence intervals around the average correlation values. The dashed lines represent the average initial correlation without added sparsity. Induced sparsity in scRNA-seq data can inflate correlations in co-expression networks, highlighting the need for alternative scRNA-seq data representations.
III.4 Results on the Simulated Data
@@
To evaluate performance of the clustering methods on datasets with known ground-truth, we employ five widely used measures to compute coherence between the identified clusters and the ground-truth clusters: Adjusted Rand Index (ARI), Normalized Mutual Information (NMI), Adjusted Mutual Information (AMI), Clustering Accuracy (ACC), and F1 Score. All evaluation measures are invariant to permutations, with a score of 0 representing random labeling and a score of 1 indicating identical clusters. K-means is used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
Figure 4: Comparing Performances of Clustering Methods on Simulated scRNA-seq Data Across Modularity. The performance of clustering methods is measured by ARI, NMI, ACC, AMI, and F1 across different levels of data modularity. The x-axis represents parameters determining data modularity, including the average number of co-expressed genes per cell type and the number of embedded modules. Our proposed methods, DIPHW, CoMem, and CoMem-DIPHW, are highlighted with dashed lines. CoMem-DIPHW, in particular, maintains strong performance and is least affected by decreases in modularity, performing significantly better even under conditions of extremely weak modularity. This behavior is observed consistently across all evaluation measures. Experiments are repeated 10 times for each parameter setting, and the 95% confidence intervals (CIs) are shown. The bar plots represent Barber’s bipartite modularity, normalized so that the most modular graphs have modularity = 1. K-means is used to cluster the output of all embedding-based methods that do not include an inherent clustering mechanism.
@@
We evaluate the performance of clustering methods under varying data modularity conditions. Modularity is controlled by two key parameters: the average number of co-expressed genes per cell type and the number of embedded cell modules. The number of modules corresponds to the number of cell types in the simulated scRNA-seq data. Weak modularity refers to a small number of co-expressed genes per module or a large number of embedded modules, which makes data structures more difficult to detect. The modularity bar plots in Fig. 4 and the heatmap visualization in Fig. S11 show that the simulated scRNA-seq data exhibit stronger modularity when the average number of co-expressed genes per cell type is higher or when the number of modules is smaller. We implemented Barber’s bipartite modularity [barber2007modularity], as shown in the bar plots in Fig. 4, to quantify the modularity of the underlying data for clustering and to analyze the association between modularity and clustering performance. Clustering performance generally improves as data modularity increases. Our proposed methods, DIPHW, CoMem, and CoMem-DIPHW, demonstrate strong and highly competitive performance across varying modularity. Specifically, CoMem-DIPHW consistently remains competitive with the best-performing methods evaluated, with its advantage particularly pronounced under weak data modularity conditions. For clarity, we present the most contrasting cases from our analysis: (i) an average of 10 and 100 co-expressed genes per module in Fig. 5 and (ii) 10 and 100 embedded modules in Fig. 6.
Figure 5: Impact of Module Size on Clustering Performance by ARI. The bar plots show the mean ARI (top) and average ranking (bottom) for each method, ordered by their average ARI. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest ARI values or best ARI rankings. Our proposed hypergraph-based methods, highlighted in red and yellow, show significant advantages when data modularity is weak, that is, when the average number of co-expressed genes per cell type module is low (10 genes per module). Under strong modularity conditions, where the average number of genes per cell type module is high (100 genes per module), many methods (e.g., PCA, CoMem-DIPHW, CoMem, and Node2Vec) perform well and achieve average ARI scores above 0.9. These observations are consistent across NMI, ACC, AMI, and F1, as shown in Supplementary Section S6. K-means is used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
Figure 6: Impact of Module Count on Clustering Performance by ARI. The bar plots show the mean ARI (top) and average ranking (bottom) of each method, ordered by their average ARI. Our proposed hypergraph-based methods, highlighted in red and yellow, demonstrate significant advantages in weak modularity regimes, where the number of embedded modules (representing the number of cell types) in the simulated data is high (100 modules). Each experiment is repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest ARI values or best ARI rankings. These observations are consistent across NMI, ACC, AMI, and F1, as shown in Supplementary Section S6. K-means is used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
@@
In Figures 5 and 6, we compare the performance of our proposed hypergraph-based clustering methods (highlighted in red and yellow) with five embedding-based approaches, four community detection algorithms, and four state-of-the-art scRNA-seq clustering methods. K-means is used to cluster the output of all embedding-based methods that do not directly assign cluster membership. In weak modularity regimes (average number of co-expressed genes per module = 10 or number of modules = 100), CoMem-DIPHW and DIPHW show a clear performance advantage. For example, when the average number of co-expressed genes is 10, CoMem-DIPHW achieves an ARI of 0.30 and DIPHW achieves an ARI of 0.25, compared to 0.05 for the best baseline method (PCA), representing a performance improvement of nearly 450%. Under strong modularity conditions (average number of genes per module = 100 or number of modules = 10), multiple methods (PCA, CoMem-DIPHW, CoMem, and Node2Vec) performed well, achieving average ARI scores above 0.9.
@@
Additional results from intermediate parameter settings (average number of co-expressed genes per module: 20, 30, 40, 70, 80, 90; number of embedded modules: 20, 30, 40, 70, 80, 90), as measured by NMI, ACC, AMI, and F1, support the same observation. Across the 16 parameter settings under different modularity conditions, CoMem-DIPHW ranks first in all 16 settings by ACC and F1, 15 by ARI, and 10 each by NMI and AMI. DIPHW and CoMem also perform strongly. Either DIPHW or CoMem rank among the top three methods in at least 10 of 16 settings across all evaluation measures. Our hypergraph-based methods show significant advantages in weak modularity regimes, where data structures are harder to detect. Under comparatively strong modularity conditions, several methods (PCA, CoMem-DIPHW, CoMem, Node2Vec, DIPHW, and Leiden) perform well, with average NMI scores above 0.9. Notably, PCA slightly outperforms CoMem-DIPHW by ARI, NMI, and AMI under strong modularity, with differences in scores ranging from 0.01 to 0.02. These additional results are in Supplementary Section S6.
III.5 Results on the scMixology Benchmark Datasets
@@
A major challenge in scRNA-seq analysis is the lack of reliable ground truth for cell type annotation, which complicates the evaluation of clustering algorithms. In most scRNA-seq datasets, cell type labels are assigned using a combination of clustering results, differential expression (DE) analysis, and canonical marker genes. First, cells are grouped into clusters based on transcriptomic similarity using clustering algorithms such as K-means. Next, DE analysis identifies genes that are differentially expressed across clusters. Finally, these clusters are annotated based on known canonical marker genes associated with specific cell types. Because this annotation process relies on clustering results, using these labels to evaluate clustering performance introduces circular logic and bias toward the method originally used for annotation. To address this issue, we use the mixture control benchmarking dataset scMixology [tian2019benchmarking], where ground truth is established based on cell line identity.
@@
We evaluate the performance of our proposed method, CoMem-DIPHW, using two benchmarking datasets from scMixology [tian2019benchmarking]. These datasets provide ground truth labels for cell type annotation based on human lung adenocarcinoma cell lines. As shown in Table 4, tsImpute achieved the highest scores across all five evaluation measures (ARI, NMI, AMI, ACC, and F1) on the 3-class dataset, with CoMem-DIPHW, Louvain, and Leiden tied for second. On the 5-class dataset, CoMem-DIPHW achieved the best performance across all evaluation measures, followed by tsImpute and graph-sc. This result is consistent with the pattern observed in our simulation analysis (see Section III.4), where CoMem-DIPHW maintains strong performance under low modularity conditions.
Traditional Embedding and Community Detection Methods